What was johns age 4 years ago if he will be z years old in 5 years
Answer by macston(5194)

.
Current age=z-5 years
.
Age 4 years ago=current age-4 years
Age 4 years ago=(z-5 years)-4 years
Age 4 years ago=z-9 years
.